Determine the pH of a 0.4 M NaF solution at 25°C. The Ka of HF is 3.5x10-5.
Answer #1

use:

Kb = Kw/Ka

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Kb = (1.0*10^-14)/Ka

Kb = (1.0*10^-14)/3.5*10^-5

Kb = 2.857*10^-10

F- dissociates as

F- + H2O -----> HF + OH-

0.4 0 0

0.4-x x x

Kb = [HF][OH-]/[F-]

Kb = x*x/(c-x)

Assuming x can be ignored as compared to c

So, above expression becomes

Kb = x*x/(c)

so, x = sqrt (Kb*c)

x = sqrt ((2.857*10^-10)*0.4) = 1.069*10^-5

since c is much greater than x, our assumption is correct

so, x = 1.069*10^-5 M

use:

pOH = -log [OH-]

= -log (1.069*10^-5)

= 4.971

use:

PH = 14 - pOH

= 14 - 4.971

= 9.029

Answer: 9.03